OpenStudy (anonymous):

The hypotenuse of a right triangle is 20 centimeters, and the shorter leg is 12 centimeters. Find the length of the other leg. 2 square rout 2 cm 4 square rout cm 16 cm

OpenStudy (ivettef365):

use Pythagorean theorem

OpenStudy (ivettef365):

a^2 + b^2 = c^2 c = hypotenuse

OpenStudy (anonymous):

So I substitute them and get 16 cm?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

ok thanks a bunch!

OpenStudy (ivettef365):

12^2 + b^2 = 20^2 144 + b^2 = 400 b^2 = 400 - 144 b^2 = 256 b = \(\sqrt{256}\) b = 16

OpenStudy (ivettef365):

yw :)